OpenStudy (anonymous):

y = 4/x What is an equation for the function?

hero (hero):

Functions are always written with f(x) notation. y = f(x) so `f(x) = 4/x`

OpenStudy (anonymous):

this is the questions This graph of a function is a translation of y = 4/x . What is an equation for the function? this is the graph

hero (hero):

You didn't post "translation" the first time

OpenStudy (anonymous):

Sorry! can you help still?

hero (hero):

Hint: The graph moves 5 units to the left and 2 units upward

OpenStudy (anonymous):

so it would be y = 4/(x+6) + 2?

OpenStudy (anonymous):

these are the answers I have A. y = 4/(x+2) + 6 B. y = 4/(x+2) - 6 C. y = 4/(x+6) + 2 D y = 4/(x+6) - 2

hero (hero):

Yes, sorry, y = 4(x + 6) + 2 is correct

OpenStudy (anonymous):

Thanks for your help :)!!

hero (hero):

You should try graphing them yourself next time.